# Binary Stars Approaching Supermassive Black Holes: Hydrodynamics of Stellar Collisions, Mass Fallback and Partial TDEs

### v1

@@ 4.2 Parameter Space and Comments on the Results of Paper I
To determine the range of outcomes of stellar collisions in binary-SMBH encounters, we show in Figure 13 the probability distributions of the relative orbital angular momentum $L$ (left panel) and impact parameter $b_{12}/r_{12}$ (right panel) of the two colliding stars for both gentle and deep encounters, based on the simulation data of Paper I. We see that most collisions are grazing collisions, characterized by relatively large $b_{12}/r_{12}$. The hydrodynamical simulations reported in this paper show that for such grazing collisions, the outcomes differ between the encounter types: for gentle encounters, multiple collisions eventually lead to a single merger remnant; for deep encounters, the most common outcome is not a single merger remnant but rather two stars that have undergone significant dynamical perturbations.
@@
This work made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.

### v2

@@ 5.2 Parameter Space and Comments on the Results of Paper I
To determine the range of outcomes of stellar collisions in binary-SMBH encounters, we show in Figure 16 the probability distributions of the relative orbital angular momentum $L$ (left panel) and impact parameter $b_{12}/r_{12}$ (right panel) of the two colliding stars for both gentle and deep encounters, based on the simulation data of Paper I. We see that most collisions are grazing collisions, characterized by relatively large $b_{12}/r_{12}$. The hydrodynamical simulations reported in this paper show that for such grazing collisions, the outcomes differ between the encounter types: for gentle encounters, multiple collisions eventually lead to a single merger remnant; for deep encounters, the outcome is highly sensitive to the collision parameters and initial stellar structure. In our representative deep-grazing encounters, the polytropic case (Run 2) produces two perturbed remnants, whereas the MESA case (Run 6) produces a single merger. Because the contact conditions are not identical—see Table 1—we refrain from attributing the different outcomes to stellar structure alone. At contact, Run 2 has $v_{12}\approx 1.354v_{\rm esc}$ and $b_{12}/r_{12}\approx 0.489$, whereas Run 6 has $v_{12}\approx 1.012v_{\rm esc}$ and $b_{12}/r_{12}\approx 0.525$. Hence the contact kinetic energy in Run 2 is $\sim(1.354/1.012)^{2}\approx 1.8\times$ that of Run 6, and the product $v_{12}(b_{12}/r_{12})$ is $\sim 25\%$ larger—both trends disfavoring capture/merger in Run 2. Small differences in contact speed and geometry can shift the balance among orbital binding, tidal energy deposition, and mass stripping; our results therefore illustrate the possible outcomes under these specific initial conditions, and a controlled survey varying one factor at a time will be needed to isolate the role of stellar structure. This interpretation is consistent with the classical tidal-capture scalings in which the dissipated energy rises steeply with decreasing pericenter distance (e.g., Press & Teukolsky, 1977; Lee & Ostriker, 1986).
@@
We thank the anonymous referee for suggestions that have improved the manuscript. This work made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University. The computations were supported by the National Science Foundation of China (grant No. 12173024).